### v1

@@ V Examples of absorption enhancement and the bistability
In examples we consider a spherical silver nanoparticle with a polymer shell in the water. The refractive index of most polymers is close to the glass refractive index $1.55$ Zhang et al. (2020). The water refractive index is $1.33$. For simplicity, we take the same refractive index $\eta=1.33$ for the nanoparticle shell and the water environment.
@@
Nanoparticle resonant polarizability is calculated in Appendix taking into account the radiative losses Bottcher (1973); Protsenko and Uskov (2012) and the losses due to collisions of the metal electrons with the nanoparticle surface Kik and Brongersma (2007); Protsenko and Uskov (2012). We estimate the LPR wavelength $\lambda_{LPR}=408$ nm and the half-wigth $\Gamma\approx 80$ THz. We take the nanoparticle radius $r_{p}=13$ nm corresponding to the maximum LPR quality factor $Q=29$. Resonant absorbers are on the nanoparticle dielectric shell of the thickness $\Delta r\geq 5$ nm. We take the minimum shell thickness $5$ nm because of a rapid de-population of the absorber excited states (fluorescence quenching) at $\Delta r<5$ nm due to the absorber interaction with the high-order (quadrupole etc.) plasmonic modes Dulkeith et al. (2005). The number of absorbers on the shell is large $N_{m}\geq 10$.
@@
We consider the stationary population $n_{1}$ of the upper level of the absorber resonant transition as a characteristic of the absorption. We compare $n_{1}$ with and without the nano-particle, so we examine the LPR-stimulated absorption enhancement (or suppression) at various numbers of absorbers $N_{m}$, the dielectric shell thickness $\Delta r$, the absorber upper-level de-population time $\tau$ and the absorber transition half-width $\gamma$.
[…]
@@
Figure 3: (a) The stationary upper-level population per absorber versus the normalized external field intensity at various number of absorbers $N_{m}=10$ (curve 1), $80$ (2), and $120$ (3). Curve 4 is for the absorber upper-level population in free space. (b) The ratio $R$ of the emitter upper-level population with and without the nanoparticle (the LPR absorption enhancement) for the same parameters as in Figure 3a. The unstable stationary absorption corresponds to dashed parts of curves 3.
@@
Curves 3 in Figs. 3a, b are for $N_{m}=10$, when $\alpha_{p}/\alpha_{p}^{max}\approx 1.8\sim 1$ and both conditions (28) and (29) (or (30)) are satisfied. We see from curves 1 on Figs. 3 that the LPR absorption is enhanced, relative to the free space absorption, at any input field intensity. The maximum enhancement is about 2.5 –3 times, respectively to free space, at a small field intensity far from the absorber transition saturation.
@@
We increase the number of absorbers. Curves 2 in Figs. 3 are for $N_{m}=80$ so $\alpha_{p}/\alpha_{p}^{max}=14.4\gg 1$ and the condition (28) for a small collective emission and the absorber-to-particle energy transfer is not satisfied. We see the absorption suppression by LPR at small $\sigma_{m}P\tau<1.6$. Otherwise, there is the absorption enhancement for a large $\sigma_{m}P\tau$, when the collective spontaneous emission and the energy transfer are reduced due to the absorber transition saturation. Note that $A=0.97<A_{cr}$ for curves 2, so there is no absorption bistability.
@@
We increase $N_{m}$ more and take $N_{m}=120$ for curves 3 in Figs. 3, so $A=11.6>A_{cr}=8$, the condition (32) is satisfied and we see the bistability in curves 3. The bistability is between regimes with the low and the high absorption. The LPR suppresses the absorption at a small and enhances it at a large external field intensity. The unstable stationary solution corresponds to the dashed parts of curves 3.
@@
Figure 4: $n_{1}$ (a) and $R$ (b) versus the normalized input field intensity for $N_{m}=120$ and different thickness of the nanoparticle shell $\Delta r=5$ nm (curves 1), $6$ nm (2) and 9 nm (3). The rest of the parameters is the same as for Figure 3. The absorber without the nanoparticle corresponds to curve 4. The dashed parts of curves mean the unstable stationary solution. The bistability disappears with the increase of the shell thickness $\Delta r$.
@@
Figure 4 shows $n_{1}$ and $R$ dependencies on the normalized external field intensity $\sigma_{m}P\tau$ for a large number of absorbers $N_{m}=120$ and different thicknesses of the nanoparticle shell $\Delta r=5$ (curves 1), $6$ (2) and $9$ nm (3). We see that the bistability disappears with only 1 nm increase of the nanoparticle shell thickness. For each shell thickness, the absorption is suppressed at small input field intensities and enhanced at high intensities, when the collective emission and the absorber-to-particle energy transfer are suppressed due to the absorber transition saturation.
@@
Let us see what happens with the decrease of the absorber state de-population rate (i.e. the increase $\tau$).
@@
Figure 5: $n_{1}$ (a) and $R$ (b) versus the normalized input field intensity for $N_{m}=10$ and different absorber upper level de-population time $\tau=2$ ps (curves 1), $20$ ps (2) and 50 ps (3). Curve 4 is for an absorber without the nanoparticle. The shell thickness $\Delta r=5$ nm, the rest of parameters is the same as for Figure 3. The upper-level population and the absorption are reduced with the increase of $\tau$.
@@
The absorption enhancement by LPR is reduced with the increase of $\tau$, from curves 1 to curves 3 in Fig 5, because of the growing of the spontaneous emission to the LPR contribution to the population decay.
@@
Suppose we have a large absorber upper state de-population time $\tau=1/2\gamma_{sp}=0.5$ ns; such a de-population is, for example, because of the spontaneous emission to free space with the typical optical domain rate $\gamma_{sp}=1$ GHz Scully and Zubairy (1997). We assume the absorber transition half-width $\gamma=100$ GHz. Then $\alpha_{p}/\alpha_{m}=20<2\gamma\tau=100$ so the condition (29) is not satisfied and there is no absorption enhancement by the LPR at any external field intensity. Meanwhile, the bistability in absorption exists, as shown by curve 3 in Figure 6, when the bistability condition (32) is true.
@@
Figure 6: The upper absorber state population $n_{1}(\sigma_{m}P\tau)$ at the suppression of absorption by the spontaneous emission to LPR mode, when $\tau=0.5$ ns and $\gamma=100$ GHz. The number of absorbers $N_{m}=10$ is for curve 1; $N_{m}=16$ is for curve 2 and $N_{m}=30$ is for curve 3; curve 4 is for the absorber in free space. The bistability condition (32) is satisfied and the bistability in absorption exists for curve 3. The dashed part of curve 3 corresponds to the unstable stationary state.
VI Discussion
@@
An electric field enhancement near plasmonic nanoparticles increases the resonant absorption of atoms, molecules, semiconductor media, etc. The estimation (27) shows the growth of the absorption cross-section near plasmonic particle in a factor $\alpha_{p}^{2}/4r^{6}=9.5$ – without taking into account the spontaneous emission of the absorber to the LPR mode (described by the denominator in Eq. (27)). Spontaneous and collective emissions to the LPR mode significantly reduce the absorption making it, at some conditions, smaller than in the free space. In our examples, the maximum absorption enhancement near plasmonic nanoparticles is about the factor 3 (see curve 1 in Figure 3b), with the spontaneous emission to the LPR mode taken into account.
@@
The maximum absorption enhancement is found for a small external field when the absorbing transition is not saturated. A noticeable absorption enhancement is possible at a fast (on ps time scale) de-population of the upper absorbing state, when the rate of spontaneous emission to the LPR mode is about or less than the de-population rate due to other processes as, for example, the decay of absorber excitation to the non-resonant high-order plasmonic modes Dulkeith et al. (2005). When the population decay is not so fast, the relative suppression of the absorber excitation by the radiation to the LPR is large and we do not have the absorption enhancement, as shown in Figure 6. The condition (29) must be satisfied for the LPR absorption enhancement. Spontaneous and collective emissions to the LPR mode are important factors, they can not be neglected in the resonant absorption near metal nanoparticles.
@@
The spontaneous emission enhancement by the cavity mode is well-known in the laser theory Oraevskii (1994) as a Pursell effect Purcell (1946). The metal nanoparticle is analogous to the cavity for the LPR mode, so such a cavity enhances the spontaneous emission similar to a laser cavity.
@@
When a few resonant absorbers are placed near metal nanoparticles, a collective emission from absorbers to the plasmonic mode is possible. A photon emitted by the atom $i$ is re-absorbed by the atom $j$ and emitted again by the collective emission. This is a Dicke effect Gross and Haroche (1982) with the probability (and, therefore, the rate) larger than for the photon emission by a single atom. The collective emission de-populates the absorber’s upper state. Collective emission to the LPR mode is similar to the collective emission to a lasing mode Leymann et al. (2015); Jahnke et al. (2016); André et al. (2019); Protsenko and Uskov (2022). The collective emission is a nonlinear process, its rate depends on the emitting state populations. In particular, the collective emission is suppressed at the saturation of the emitting transition.
@@
The particle-absorber near-field interaction is a mutual process: the metal nano-particle LPR mode excites the polarization and transfers the energy to the absorber through the near field and vice-versa: absorbers excite the particle polarisation through the near field. Such an energy transfer is the origin of the suppression factor in the denominator in Eq. (14) for the absorber polarisation. The collective emission and the emitter-particle interaction depend on the absorber states population. Such a dependence leads to a positive feedback: the absorption is increased with the upper-level population at certain conditions. The feedback is fast: note the square in the nonlinear denominator in Eq. (26). So, at certain parameter values, the feedback nonlinearity "overtakes" the transition saturation and leads to a decrease of the upper-level population $n_{1}$ with the external field intensity. When $n_{1}$ decreases, the stationary absorption is unstable, see the dashed parts of curves in Figs. 3, 4, and 6. Such anomalous absorption causes absorption bistability. The bistability exists at the condition (32) which can be satisfied at the enhancement (when condition (29) is true) or the suppression (when condition (29) is not true) of absorption by LPR, see Figures 4 and 6. With the LPR absorption enhancement, the bistability is between the absorption suppression and the absorption enhancement regimes, see Figures 3 and 4.
@@
The optical bistability is possible in plasmonic nanolaser with external signal Protsenko et al. (2008). The model of Protsenko et al. (2008) involves a single emitter near the plasmonic nanoparticle and does not include the spontaneous and collective emissions to the LPR mode.
@@
The physical nature of a well-known absorptive optical bistability in a cavity with a two-level absorber (or absorbers) Koch, Schmidt, and Haug (1985) is the resonant transition saturation. This is different from the physics of the bistability in the resonant absorption in the LPR near metal nanoparticle described here. The absorptive bistability Koch, Schmidt, and Haug (1985) is not related to the collective emission and the near-field interaction of absorbers with the metal nanoparticle.
@@
The nanoparticle with resonant emitters is, probably, one of the smallest systems with optical bistability with the size much smaller than the optical wavelength.
@@
Here we carry out the semi-classical analysis of the resonant absorption near a metal nanoparticle. Only the term of the spontaneous emission into the LPR mode in Eq. (17) is introduced according to quantum mechanical relation (16). We note that the term of the spontaneous emission to the lasing mode appears in the laser theory the same way P. R. Rice  and Carmichael (1994); André et al. (2019). More detailed theory of the resonant absorption near plasmonic nanoparticles must include quantum fluctuations of the field and the absorber states populations. Such a quantum analysis is necessary, taking into account a small volume of the plasmonic mode Maier (2006), which keeps only one or a few photons. Such an analysis will be carried out in the future.
VII Conclusion

### v2

@@ 5 Examples of absorption enhancement and the bistability
As an example, consider a spherical silver nanoparticle with a polymer shell in the water. The refractive index of most polymers is close to the refractive index of glass $1.55$ [54]. The refractive index of water is $1.33$. For simplicity, we assume the same refractive index $\eta=1.33$ for the nanoparticle shell and the water environment.
@@
The resonant polarisability of the nanoparticles is calculated in the Appendix, taking into account the radiative losses [55, 35] and the losses due to collisions of the metal electrons with the nanoparticle surface [56, 35]. We estimate the LPR wavelength $\lambda_{LPR}=408$ nm and the half width $\Gamma\approx 80$ THz. We take the nanoparticle radius $r_{p}=13$ nm, which corresponds to the maximum LPR quality factor $Q=29$. The resonant absorbers are located on the nanoparticle dielectric shell of the thickness $\Delta r\geq 5$ nm. We chose a minimum shell thickness, because of a rapid increase in absorber fluorescence quenching near metal nanoparticles has been observed at smaller $\Delta r$ [34, 57]. Strong quenching at small $\Delta r$ can be caused by several mechanisms, besides the dipole-dipole interaction, such as a nanometal surface energy transfer (NSET) [58]. Taking $\Delta r\geq 5$ nm we can neglect by the nanoparticle-absorber’s interactions, but the dipole-dipole one, to keep our system as simple as possible.
The number of absorbers on the shell is large $N_{m}\geq 10$.
@@
We consider the stationary population $n_{1}$ of the upper level of the absorber resonance transition as a property of the absorption. We compare $n_{1}$ with and without the nanoparticle, so we examine the LPR-stimulated absorption enhancement (or suppression) at different numbers of absorbers $N_{m}$, the dielectric shell thickness $\Delta r$, the absorber upper level depopulation time $\tau$ and the absorber transition half width $\gamma$.
Suppose we have a fast depopulation of the upper absorber states with a population decay time $\tau=2$ ps. Such a depopulation determines the absorber transition half-width so $\gamma=1/2\tau$. We take the parameter values written above and find that the condition (29) for the LPR absorption enhancement is satisfied: $\alpha_{p}/\alpha_{m}=100>2\gamma\tau=1$. Figure 3a shows $n_{1}$, Figure 3b - the ratio $R=n_{1}/n_{1}^{(0)}$ at different $N_{m}$, as a function of the normalised external field intensity $\sigma_{m}P\tau$; $n_{1}^{(0)}$ is an absorber upper level population calculated without the nanoparticle. At current parameter values, $\sigma_{m}P\tau=1$ corresponds to an external field intensity 0.54 MWt/cm2. Such intensities can be achieved, for example, in a pulsed laser regime with pulse duration $\sim 1$ ns $\gg\tau=2$ ps, so that we have a quasi-stationary population of absorber states. The external field energy is about 0.5 mJ/pulse at a pulse cross section of 1 cm2.
@@
Figure 3: (a) The steady-state upper-level population per absorber versus the normalised external field intensity for different numbers of absorbers $N_{m}=10$ (curve 1), $80$ (2), and $120$ (3). Curve 4 is for the absorber upper-level population in free space. (b) The ratio $R$ of the absorber upper-level population with and without the nanoparticle (the LPR absorption enhancement) for the same parameters as in Figure 3a. The unstable steady-state absorption corresponds to the dashed parts of curves 3.
@@
Curves 3 in Figs. 3a, b are for $N_{m}=10$, when $\alpha_{p}/\alpha_{p}^{max}\approx 1.8\sim 1$ and both conditions (28) and (29) (or (30)) are satisfied. We can see from curves 1 in Figs. 3 that the LPR absorption is enhanced, relative to the free space absorption, at any input field intensity. The maximum enhancement is about 2.5 –3 times that of free space at a small field intensity far from the absorber transition saturation.
@@
We increase the number of absorbers. Curves 2 in Figs. 3 are for $N_{m}=80$ so $\alpha_{p}/\alpha_{p}^{max}=14.4\gg 1$ and the condition (28) for a small collective emission and energy transfer from absorber to particle is not satisfied. We see the absorption suppression by LPR at small $\sigma_{m}P\tau<1.6$. Otherwise, there is the absorption enhancement for a large $\sigma_{m}P\tau$, when the collective spontaneous emission and the energy transfer are reduced due to the saturation of the absorber transition. Note that $A=0.97<A_{cr}$ for curves 2, so there is no absorption bistability.
@@
We increase $N_{m}$ further and take $N_{m}=120$ for curves 3 in Figs. 3, so $A=11.6>A_{cr}=8$, the condition (32) is satisfied and we see the bistability in curves 3. The bistability is between regimes with the low and the high absorption regimes. The LPR suppresses the absorption at low external field intensities and enhances it at high external field intensities. The unstable steady state solution corresponds to the dashed parts of curves 3.
@@
Figure 4: $n_{1}$ (a) and $R$ (b) versus the normalised input field intensity for $N_{m}=120$ and different nanoparticle shell thickness $\Delta r=5$ nm (curves 1), $6$ nm (2) and 9 nm (3). The rest of the parameters are the same as in Figure 3. The absorber without the nanoparticle corresponds to curve 4. The dashed parts of the curves represent the unstable steady-state solution. The bistability disappears with the increase of the shell thickness $\Delta r$.
@@
Figure 4 shows the dependence of $n_{1}$ and $R$ on the normalised external field intensity $\sigma_{m}P\tau$ for a large number of absorbers $N_{m}=120$ and different nanoparticle shell thicknesses $\Delta r=5$ (curves 1), $6$ (2) and $9$ nm (3). We see that the bistability disappears with only 1 nm increase in the nanoparticle shell thickness. For each shell thickness, the absorption is suppressed at low input field intensities and enhanced at high intensities, when the collective emission and the absorber-to-particle energy transfer are suppressed due to the absorber transition saturation.
@@
Let us see what happens with the decrease of the depopulation rate of the absorber state (i.e. the increase of $\tau$).
@@
Figure 5: $n_{1}$ (a) and $R$ (b) versus the normalised input field intensity for $N_{m}=10$ and different absorber upper level depopulation time $\tau=2$ ps (curves 1), $20$ ps (2) and 50 ps (3). Curve 4 is for an absorber without the nanoparticle. The shell thickness $\Delta r=5$ nm, the rest of the parameters are the same as in Figure 3. The upper level population and the absorption are reduced with the increase of $\tau$.
@@
We first consider the case without the bistability and with a large absorption enhancement, as shown in Fig. 3 by the curve 1 for $N_{m}=10$ and $\tau=2$ ps. The absorption enhancement by LPR is decreases with increasing $\tau$, from curves 1 to curves 3 in Fig 5, due to the increase of the spontaneous emission to the LPR contribution to the population decay.
Let us now increase the absorber state depopulation time $\tau$ at a large number of absorbers $N_{m}=120$, when the condition (32) is satisfied and we have the absorption bistability. We see in Fig. 6 that the bistability region grows with the increase of $\tau$, while the absorption enhancement is very small.
Figure 6: The upper absorber state population $n_{1}(\sigma_{m}P\tau)$ at the absorption bistability, when $N_{m}=120$ and different $\tau=2$ ps (curve 1), $20$ ps (curve 2) and $50$ ps (curve 3). Unstable parts of the curves are indicated by the dashed lines. The bistability region increases as $\tau$ increases. Curve 4 is for the absorber in free space. The absorption enhancement (i.e. the positive difference in the upper parts of curves 1-3 and 4) is quite small.
@@
Suppose we have a large absorber upper state depopulation time $\tau=1/2\gamma_{sp}=0.5$ ns; such depopulation occurs, for example, due to spontaneous emission into free space with the typical optical domain rate $\gamma_{sp}=1$ GHz [52]. We assume that the absorber transition has a half-width $\gamma=100$ GHz. Then $\alpha_{p}/\alpha_{m}=20<2\gamma\tau=100$ so the condition (29) is not satisfied and there is no absorption enhancement by the LPR at any external field intensity. Meanwhile, there is the bistability in the absorption, as shown by curve 3 in Figure 7, when the bistability condition (32) is satisfied.
@@
Figure 7: The upper absorber state population $n_{1}(\sigma_{m}P\tau)$ with suppression of absorption by the spontaneous emission in the LPR mode, when $\tau=0.5$ ns and $\gamma=100$ GHz. The number of absorbers $N_{m}=10$ is for curve 1; $N_{m}=16$ is for curve 2 and $N_{m}=30$ is for curve 3; curve 4 is for the absorber in free space. The bistability condition (32) is satisfied and there is bistability in absorption exists for curve 3. The dashed part of curve 3 corresponds to the unstable steady state. There is no the absorption enhancement.
6 Discussion
@@
An electric field enhancement near plasmonic nanoparticles increases the resonant absorption of atoms, molecules, semiconductor media, etc. The estimation (27) shows the increase of the absorption cross section near plasmonic particle by a factor $\alpha_{p}^{2}/4r^{6}=9.5$ – without taking into account the spontaneous emission of the absorber to the LPR mode (described by the denominator in Eq. (27)). Spontaneous and collective emissions to the LPR mode significantly reduce the absorption making it smaller than in the free space under some conditions. In our examples, the maximum absorption enhancement near plasmonic nanoparticles is about a factor of 3 (see curve 1 in Figure 3b), taking into account the spontaneous emission to the LPR mode.
@@
The maximum absorption enhancement is found for a small external field when the absorbing transition is not saturated. A noticeable absorption enhancement is possible with a fast (on ps time scale) depopulation of the upper absorbing state, when the rate of spontaneous emission to the LPR mode is close to or less than the depopulation rate due to other processes. When the population decay is not so fast, the relative suppression of the absorber excitation by the radiation to the LPR is large and we do not have the absorption enhancement, as shown in Figure 7. For LPR absorption enhancement, condition (29) must be satisfied. Spontaneous and collective emissions to the LPR mode are important factors that cannot be neglected in the resonant absorption near metal nanoparticles.
@@
The enhancement of spontaneous emission by the cavity mode is well known in the laser theory [59] as the Pursell effect [60]. The metal nanoparticle is analogous to the cavity for the LPR mode, so such a cavity will enhance the spontaneous emission similar to a laser cavity.
@@
When a few resonant absorbers are placed close to metal nanoparticles, collective emission from the absorbers into the plasmonic mode is possible. A photon emitted by the atom $i$ is re-absorbed by the atom $j$ and re-emitted by the collective emission. This is a Dicke effect [61] with the probability (and hence the rate) is greater than for the photon emission from a single atom. The collective emission depopulates the upper state of the absorber. Collective emission to the LPR mode is similar to the collective emission to a lasing mode [62, 63, 28, 29]. Collective emission is a non-linear process, its rate depends on the populations of the emitting states. In particular, collective emission is suppressed at the saturation of the emitting transition.
@@
The particle-absorber near-field interaction is a two-way process: the LPR mode of the metal nano-particle excites the polarisation and transfers the energy through the near-field of the absorber, and vice versa: absorbers excite the particle polarisation through the near-field. Such an energy transfer is the origin of the suppression factor in the denominator in Eq. (14) for the absorber polarisation. The collective emission and the absorber-particle interaction depend on the absorber state population. Such a dependence leads to a positive feedback: the absorption is increased with the upper level population under certain conditions. The feedback is fast: note the square in the non-linear denominator in Eq. (26). Thus, for certain parameter values, the feedback nonlinearity ”overtakes” the transition saturation and leads to a decrease of the upper-level population $n_{1}$ with the external field intensity. As $n_{1}$ decreases, the steady-state absorption is unstable, see the dashed parts of the curves in Figs. 3, 4, and 6. Such anomalous absorption causes absorption bistability. The bistability exists under the condition (32), which can be satisfied by the enhancement (when condition (29) is true) or the suppression (when condition (29) is not true) of absorption by LPR, see Figures 4 and 6. In the case of LPR absorption enhancement, the bistability is between the absorption suppression and absorption enhancement regimes, see Figures 3 and 4.
@@
Optical bistability is possible in plasmonic nanolasers with an external signal [26]. The model of [26] includes a single emitter near the plasmonic nanoparticle and does not consider the spontaneous and collective emissions to the LPR mode.
@@
The physical nature of the well-known absorptive optical bistability in a cavity with a two-level absorber (or absorbers) [64] is the saturation of the resonant transition. This is different from the physics of the resonant absorption bistability in the LPR near metal nanoparticles described here. The absorptive bistability [64] is not related to the collective emission and the near-field interaction of the absorbers with the metal nanoparticle.
@@
The nanoparticle with the nearby resonant absorbers (molecules) is probably one of the smallest systems with optical bistability, with a size much smaller than the optical wavelength.
We have shown that the absorption enhancement near the metal nanoparticle and the bistability are strongly influenced by nonlinear collective absorption of the nanoparticle and nearby molecules, which occurs at high light power densities ($>1~{}MW/cm^{2}$). Such power densities for single spherical metal nanoparticles can only be achieved with powerful pulsed lasers. This makes potential applications in photovoltaics, photocatalysis, sensors, etc difficult. At the same time, the necessary power densities are easily achieved in plasmonic nanosystems with strong spatial localisation of the electromagnetic field in a plasmonic gap mode even at moderate levels of external light excitation ($\sim 1~{}W/cm^{2}$ and less) [16, 17, 18, 19]. This makes the study of the collective absorption near the plasmonic gap mode to be an interesting and important task. The present approach can be directly generalised to the plasmonic gap case.
@@
Here we perform a semi-classical analysis of the resonance absorption near a metal nanoparticle. Only the term of the spontaneous emission in the LPR mode in Eq. (17) is introduced according to the quantum mechanical relation (16). We note that the term of the spontaneous emission into the laser mode appears in the same way in the laser theory [65, 28]. A more detailed theory of the resonant absorption near plasmonic nanoparticles must include quantum fluctuations of the field and the absorber state populations. Such a quantum analysis is necessary, taking into account a small volume of the plasmonic mode [66], holding only one or a few photons. Such an analysis will be performed in the future.
7 Conclusion